OpenStudy (anonymous):

Find the focus of the parabola y = x^2 + 5x

OpenStudy (anonymous):

\[y = x ^{2}+5x\]

terenzreignz (terenzreignz):

Somehow, you have to express it in the form \[\Large (x-\color{red}a)^2 =4p(y-\color{blue}b)\]

terenzreignz (terenzreignz):

We need to complete the square.... do you know how to do that? :)

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Well then, complete the square for the x-part

OpenStudy (anonymous):

\[\left( x ^{2} + 2.25\right)^{2} = y+6.25\]

OpenStudy (anonymous):

well, not x^2 in the parentheses

OpenStudy (anonymous):

What exactly does the p stand for?

terenzreignz (terenzreignz):

hang on...

terenzreignz (terenzreignz):

Where did 2.25 come from

OpenStudy (anonymous):

2.5 is what I meant

terenzreignz (terenzreignz):

Oh, okay, great :) So... \[\Large \left(x +\frac52\right)^2 = y+\frac{25}{4}\] (I used fractions, but it's essentially the same)

terenzreignz (terenzreignz):

Okay, so about that p value... that's actually the key-value here :) You need to find the value of p, so... use this equation... \[\Large \left(y+\frac{25}4\right)=4\cdot \frac14 \left(y+\frac{25}4\right)\]

terenzreignz (terenzreignz):

that part clear? ^

OpenStudy (anonymous):

Yes, that makes sense

terenzreignz (terenzreignz):

So, rewrite... \[\Large \left(x +\frac52\right)^2 =4\cdot \frac14\left( y+\frac{25}{4}\right)\]

OpenStudy (anonymous):

What do we do with the 1/4 now?

terenzreignz (terenzreignz):

That is your p-value. So, you now want to find the vertex of your parabola. Anyway, now that you have completed the square, it shouldn't be hard.... \[\Large y = \left(x+\frac52\right)^2-\frac{25}4\] What's the vertex?

OpenStudy (anonymous):

-5/2?

terenzreignz (terenzreignz):

Vertex. The whole vertex, and not just the x-value of the vertex. The vertex is the point \[\Large \left(-\frac52 \ , \ -\frac{25}4\right)\]

OpenStudy (anonymous):

After we find the vertex, how do we find the focus?

terenzreignz (terenzreignz):

Sorry, I think I was very vague when guiding you through this... Three things we need to find the focus... Vertex (check) p-value (check) concavity Which way does this parabola open?

OpenStudy (anonymous):

Upwards

terenzreignz (terenzreignz):

Right. So we now have all we need to find the focus. the FOCUS is the point that is exactly p-units away from the vertex in the UPWARD direction :)

OpenStudy (anonymous):

So , -24/4 = -6. Thank you very much! Just one more question. How do you determine if it is concave up or concave down? (I had an image of the graph that I looked at for this one).

terenzreignz (terenzreignz):

Wait, -6 is NOT the focus, that's just the y-value of the focus... just sayin' the focus is \[\Large \left(-\frac52 \ , \ -6\right)\] but other than that, good job :)

terenzreignz (terenzreignz):

Now, the concavity is determined by the coefficient of \(\large x^2\). If it's negative, then the parabola opens down. If it's positive, then the parabola opens up

OpenStudy (anonymous):

haha, thanks for keeping me straight

OpenStudy (anonymous):

Okay, great. Thanks so much!